### v2

# A Controllable and Realistic Framework for Evaluating Microservice Scheduling in Cloud-Edge Continuum

Abstract.
The transition from traditional architectures to containerized microservices within the cloud-edge computing continuum introduces significant challenges, particularly in the efficient scheduling of microservices under dynamic conditions. Complex and fluctuating call-graph dependencies, varying cross-node communication latencies, and unpredictable bandwidth conditions substantially impact the performance and reliability of deployed microservices. Consequently, accurately evaluating scheduling policies in such dynamic environments remains essential yet challenging due to the lack of realistic and controllable evaluation frameworks.
In this paper, we propose iDynamics, a novel evaluation framework designed explicitly to address these challenges. iDynamics provides comprehensive and controllable evaluation capabilities by emulating realistic dynamics, including configurable call-graph topologies, cross-node communication delays, and bandwidth variability. The framework is composed of modular components, such as the Graph Dynamics Analyzer, Networking Dynamics Manager, and Scheduling Policy Extender, enabling fine-grained environmental control and facilitating systematic comparisons of different scheduling strategies. Extensive experiments on a real cloud-edge testbed demonstrate that iDynamics effectively captures diverse dynamic scenarios encountered in microservice deployments, offering a robust solution for evaluating and optimizing policy performance under realistic and controllable conditions.
Microservices, Policy Evaluation, Controllable Dynamics, Cloud-Edge Continuum
copyright: noneconference: ACM Conference; Month YYYY; City, Countryccs: Computer systems organization Cloud computingccs: Networks Network dynamicsccs: Software and its engineering Scheduling
1. Introduction
Microservices have emerged as an important architecture in the cloud-edge computing continuum, revolutionizing the design and implementation of distributed applications. Characterized by their modularity and decentralization, microservices enhance flexibility and scalability, effectively meeting dynamic application demands (Pallewatta et al., 2023; Cerny et al., 2018; Boucher et al., 2018). However, this architectural transition introduces significant challenges, especially for managing complex workflows and intricate service dependencies (Gu et al., 2021; Luo et al., 2024; Pallewatta et al., 2024; Xu et al., 2022a). Unpredictable user requests, varying execution paths across services, and fluctuating network conditions exacerbate the difficulty of ensuring SLA (Service Level Agreement) compliance for applications deployed across cluster nodes.
Within the cloud-edge continuum, numerous microservices interact frequently and are often shared among multiple applications (Liu et al., 2023; Luo et al., 2022). Such interactions can lead to performance degradation due to resource contention and cascading delays along request execution paths, directly impacting critical SLA-related metrics such as latency and throughput.
Given the complexities arising from workload variability, dynamic application call-graphs, and network-level fluctuations, fine-grained scheduling strategies and robust policy evaluation frameworks are essential. A well-designed policy evaluation framework would enable cloud providers to more accurately and efficiently accommodate diverse microservice applications while adhering to different SLA requirements.
Although various frameworks have been proposed for evaluating scheduling policies for cloud microservices, existing solutions still have limitations. For instance, AutoARMOR (Li et al., 2021) automates inter-service access control policy generation but overlooks dynamic networking conditions and is limited in supporting diverse SLA evaluations. OptTraffic (Zhu et al., 2023) optimizes containerized microservice traffic but neglects cross-node latency fluctuations and introduces complexity due to pairwise analysis of dependent container replicas. Similarly, NetMARKS (Wojciechowski et al., 2021) utilizes dynamic network metrics from Istio Service Mesh for pod scheduling, yet it lacks comprehensive bi-directional traffic analysis and can lead to imbalanced load distributions. Other existing methods (Sun et al., 2024; Song et al., 2023; Cao and Sharma, 2021; Marchese and Tomarchio, 2022b, a) also face limitations regarding handling call-graph dynamics and networking variations.
To address these gaps, we propose iDynamics, a novel framework that systematically manages multiple layers of dynamics, enabling precise evaluation of scheduling policies under realistic conditions. With iDynamics, users can observe the impact of dynamics on application performance and identify SLA violations effectively. Figure 4 illustrates the high-level design of our proposed framework, highlighting its modular components. To our knowledge, iDynamics is the first unified evaluation framework specifically designed with controllable dynamics for microservice applications in cloud-edge continuum environments. Our primary contributions are:
  • A novel evaluation framework tailored for microservices under dynamic cloud-edge conditions.
  • Comprehensive design and implementation of modular components, including a Graph Dynamics Analyzer, Networking Dynamics Manager, and Scheduling Policy Extender.
  • Extensive validation experiments conducted on a real-world cloud-edge testbed.
The remainder of this paper is structured as follows: Section 2 reviews related work. Section 3 presents the background and motivation. Section 4 introduces the iDynamics framework and its components. Section 5 describes the design of the Graph Dynamics Analyzer, and Section 6 details the Networking Dynamics Manager. Section 7 explains the Scheduling Policy Extender. Section 8 provides performance evaluations and detailed case studies. Finally, Section 9 concludes the paper and outlines future directions.
2. Related Work
The management of microservices in different environments has been extensively explored. This section reviews existing literature with an emphasis on microservice management, call-graph analysis, network dynamics emulation, and network-aware scheduling.
2.1. Microservice Management
FIRM (Qiu et al., 2020) leverages telemetry metrics and machine learning models to fine-tune microservice management by identifying SLO violations and resource contentions, subsequently mitigating violations through reprovisioning actions. Erms (Luo et al., 2022) formulates resource scaling models specifically addressing latency goals for shared microservices with extensive call graphs. GrandSLAm (Kannan et al., 2019) predicts microservice execution stages’ completion times, dynamically batching and reordering requests based on these predictions. Nevertheless, these approaches often neglect dynamic cross-node delays and the impacts of bi-directional communication between replicated microservices.
2.2. Call-graph Dynamics Analysis
Sage (Gan et al., 2021) employs graphical Bayesian models for root cause analysis of cascading QoS violations in interactive microservices, prioritizing practicality and scalability. Tian et al. (Tian et al., 2019) develop a synthetic workload generator for DAG jobs by analyzing large-scale cluster traces. Luo et al. (Luo et al., 2021) characterize microservice call graphs from Alibaba cluster data, categorizing calling dependencies into three distinct types. Parslo (Mirhosseini et al., 2021) introduces a gradient descent method for decomposing end-to-end SLO budgets into node-specific latency targets. These solutions, however, are computationally intensive and unsuitable for handling dynamically changing request patterns efficiently.
2.3. Network Dynamics Emulation
Networking dynamics significantly influence microservice performance in cloud-edge environments. THUNDERSTORM (Liechti et al., 2019) provides tools for evaluating distributed system performance under dynamic network topologies. Kollaps (Gouveia et al., 2020) introduces a decentralized emulator to simulate dynamic network conditions using simplified topologies. SplayNet (Schmid et al., 2016) extends Splay (Leonini et al., 2009) for arbitrary topology emulation via graph analysis. However, these solutions lack efficient measurement and controllable injection mechanisms essential for accurately evaluating scheduling policies. Existing delay injection methods (Qiu et al., 2020; Wang et al., 2021; Xu et al., 2022b; Fu et al., 2021b, a) typically use static or uniform delay matrices, limiting adaptability and complicating integration into dynamic cloud-edge scenarios.
2.4. Network-aware Microservice Scheduling
Network-aware scheduling strategies such as NetMARKS (Wojciechowski et al., 2021), Marchese et al. (Marchese and Tomarchio, 2022a), and OptTraffic (Zhu et al., 2023) utilize dynamic metrics and infrastructure network conditions to schedule Kubernetes pods optimally. However, these methods insufficiently handle detailed bi-directional traffic metrics and varying cross-node communication latencies, which are critical in real-world cloud-edge deployments.
3. Background and Motivation
3.1. Background
Microservices architecture breaks applications into small, independent services that communicate over networks (Luo et al., 2021; Gan et al., 2019). In cloud-edge computing, these services can run in central cloud data centers, on edge nodes (near end-users/devices), or across both. In practice, an application might be split so that some microservices run in the cloud (for centralized tasks or global data) and others run at the edge (for local data processing or quick response), communicating over the network. This hybrid deployment improves application performance by reducing the distance data must travel, lowering user-perceived latency and offloading work from the cloud to local nodes. However, despite the promising benefits of deploying microservice applications in cloud-edge clusters, there are notable challenges when the deployment goes to a practical scenario.
Figure 1. An envisioned workflow illustrating containerized microservice execution and communication under networking dynamics in a cloud-edge continuum.
3.2. Motivation
3.2.1. Dynamic Networking Conditions
In cloud-edge scenarios, microservice applications frequently experience dynamic and heterogeneous networking conditions. Figure 1 illustrates the envisioned workflow of microservices executing across multiple nodes with fluctuating cross-node latencies and bandwidth availability. Microservices heavily depend on network communications through Remote Procedure Calls (RPCs) or REST APIs, resulting in greater sensitivity to cross-node communication variations. Consequently, fluctuating cross-node latencies and bandwidth constraints significantly affect both end-to-end response times and throughput, making it critical to accurately evaluate how scheduling policies perform under such conditions.
Figure 2. Different request types triggering distinct call-graph structures (application source (Gan et al., 2019)).
3.2.2. Variability in Call-graph Structures
Microservice requests traverse complex call-graphs, which vary significantly depending on the type of incoming requests. As depicted in Figure 2, widely-used benchmark applications such as Social Network (Gan et al., 2019) feature different request types (e.g., compose-post, read-home-timeline, read-user-timeline), each triggering unique call-graph structures. Understanding these variations is crucial, as scheduling policies must handle not only diverse call-graphs but also dynamically changing proportions of request loads. Consequently, evaluating scheduling strategies requires capturing these variations in call-graph structures and their respective workloads accurately.
3.2.3. Imbalanced Traffic Distribution
[…]
Figure 3. Imbalanced traffic loads among UM-DM pairs under high workload (5k Queries Per Second) scenarios.
Thus, accurately characterizing and evaluating scheduling policies requires a systematic framework that considers: (1) dynamic network conditions, including variable latency and bandwidth; (2) variations in call-graph structures arising from diverse request types and volumes; and (3) imbalanced traffic distribution within call-graphs, which affects SLA adherence.
Figure 4. The proposed framework iDynamics for managing different dynamics in cloud-edge continuum.
4. Framework of iDynamics
Motivated by the challenges discussed previously, we propose iDynamics, a unified evaluation framework for implementing and evaluating different scheduling policies of containerized microservices in cloud-edge computing environments. To enable robust policy evaluation, iDynamics comprehensively manages various dynamic conditions, including networking variations, changing call-graph structures, and imbalanced traffic distributions. Figure 4 illustrates the framework’s main components and their interactions during the evaluation and scheduling processes. Key components of iDynamics include:
  • Graph Dynamics Analyzer: Consisting of the UM-DM Traffic Profiler and Call-graph Builder, this component analyzes dynamic call-graphs triggered by different request types and quantifies traffic imbalances under varying workloads (Section § 5).
  • Networking Dynamics Manager: Including an Emulator and a distributed Measurer to manage and emulate realistic networking conditions such as cross-node latency and bandwidth. The Emulator generates configurable dynamics, while the distributed Measurer collects actual communication metrics via daemon agents deployed across cloud-edge nodes (Section § 6).
  • Scheduling Policy Extender: Including a Policy Customization Interface and a Utility Function Module, this component enables rapid development, customization, and evaluation of scheduling strategies. Users can implement tailored scheduling logic and evaluate their policies across diverse dynamic scenarios (Section § 7).
As depicted in Figure 4, the working procedure of the proposed iDynamics framework consists of multiple stages.
[…]
    Users submit realistic or synthetic workloads (varying request types and QPS) to the deployed microservices.
[…]
    Performance metrics are collected using a metrics analysis stack (e.g., Jaeger, Istio, Prometheus). The Graph Dynamics Analyzer builds the triggered call-graph and analyzes traffic between microservices, forwarding these insights to the Scheduling Policy Extender.
[…]
    Concurrently, the Networking Dynamics Manager measures real-time cross-node networking conditions and optionally injects realistic network dynamics into the cloud-edge environment.
[…]
    Utilizing collected dynamics data (traffic, call-graphs, latency, bandwidth), users implement and evaluate customized scheduling policies via provided interfaces and utility functions.
[…]
    SLA targets are defined by users, serving as criteria to trigger customized scheduling policies.
[…]
    SLA violations for policy evaluation are induced by either tightening SLA criteria or injecting intensified dynamics (e.g., increased latency or workload), thus simulating performance degradation. Scheduling decisions are temporarily stored in a decision queue, then filtered based on compliance requirements or migration constraints, enhancing flexibility and control.
[…]
    Finally, the filtered scheduling decisions are executed to adjust the microservice deployment according to current conditions.
With iDynamics, scheduling policies can be rapidly implemented, tested, and refined under realistic yet controllable conditions. For instance, traffic-aware and latency-aware scheduling strategies can be seamlessly developed using the dynamic metrics from the Graph Dynamics Analyzer, controlled network conditions from the Networking Dynamics Manager, and flexible interfaces provided by the Scheduling Policy Extender, as demonstrated in Figure 4.
5. Design of Graph Dynamics Analyzer
Considering the characteristics of microservice call-graphs described in Section § 3.2 Motivation, the Graph Dynamics Analyzer is designed primarily to analyze topology dynamics (variations in triggered call-graphs) and traffic dynamics (imbalanced traffic distribution) of microservice applications deployed in the cloud-edge continuum. To effectively model these dynamic call-graph structures and traffic distributions, we leverage a service mesh to implement the UM-DM Traffic Profiler and Call-graph Builder as core modules within the Graph Dynamics Analyzer, as shown in Figure  4.
5.1. Service Mesh
A service mesh facilitates real-time analysis and bidirectional monitoring of traffic between multiple upstream microservice (UM) and downstream microservice (DM) replicas. A critical requirement in developing the Graph Dynamics Analyzer component for iDynamics, as shown in Figure 4, is efficiently capturing and analyzing bidirectional traffic between dependent microservices. Collecting metrics is straightforward when each microservice has a single instance; metrics can simply be gathered from the Linux proc file system. However, multiple replicas of upstream or downstream microservices significantly complicate traffic analysis. Therefore, a service mesh is introduced to efficiently observe and manage these complex, multi-instance traffic flows.
5.1.1. Istio Service Mesh Implementation
To achieve fine-grained observation into bidirectional traffic between UM and DM microservices, including their replicas, we implemented the UM-DM Traffic Profiler using the Istio service mesh [^1] Istio is an open-source service mesh employing sidecar containers to facilitate secure communication, traffic management, and load balancing. (Istio: An open source service mesh, 2024). The service mesh operates as an additional infrastructure layer integrated into containerized microservice deployments. With Istio deployed, all microservice traffic is transparently proxied via injected sidecar containers.
In Kubernetes clusters with Istio service mesh enabled, each microservice pod includes the application container alongside an Envoy sidecar container. As illustrated in Figure 5, the service mesh architecture proxies traffic among microservice containers. In the Data Plane, orange rectangles represent the microservice containers, while green rectangles denote the sidecar containers that manage traffic and maintain dependencies between services. The green lines indicate data flows and dependency relationships among microservices. In the Control Plane, the Istio service mesh manages services and interactions via components including Pilot, Mixer and Citadel for different purposes.
Figure 5. Overview of a service mesh structure consisting of data plane and control plane.
5.1.2. Overhead Analysis
Deploying a service mesh introduces additional processing delays due to the sidecar proxy layer. However, empirical evaluations indicate these delays minimally impact overall latency performance. According to (Istio: Performance and Scalability, 2024), in Istio 1.21.2 with telemetry v2 enabled, each request passing through client-side and server-side Envoy proxies incurs a latency increase of approximately 0.18 milliseconds at the 90th percentile and 0.25 milliseconds at the 99th percentile compared to baseline performance without proxies. These results, obtained using a 1kB payload at 1000 requests per second across varying client connections (2, 4, 8, 16, 32, 64), were measured at the CNCF Community Infrastructure Lab (CNCF Community Infrastructure Lab : An on-demand infrastructure resource for CNCF developers.(2024), CIL). Thus, the added latency overhead from the service mesh is negligible and acceptable for most practical scenarios.
5.2. Call-graph Builder
The primary function of the Graph Dynamics Analyzer component is to construct real-time traffic flow graphs between deployed microservice instances. Different request types and varying QPS result in dynamic call-graph structures, generating diverse and complex topology scenarios crucial for evaluating scheduling policies against varying SLA targets.
We introduce the following terminology:
Stress Element: A Stress Element (SE) encapsulates the interaction between two dependent microservices: the Upstream Microservice (UM) and the Downstream Microservice (DM). The SE quantifies traffic stress, defined as the average of the bi-directional traffic—both sent and received—between microservice pairs over a given interval. Both directions significantly influence the application’s end-to-end performance. Mathematically, the stress of a Stress Element is represented by:
(1) $\text{stress}^{\mu}_{\sigma}(\mu^{UM},\sigma^{DM},\Delta t)=\frac{\text{Bi-directional\_traffic}(\mu^{UM},\sigma^{DM})}{2\Delta t}$
[…]
Algorithm 1 constructs the call-graphs with dependency and associated traffic load based on the defined stress elements. Initially, the algorithm retrieves all running microservices deployed within a namespace, which is used to seperate different applications. It then initializes a directed graph, with each microservice represented as a graph node. For each microservice pair, the algorithm calculates the stress using real-time traffic metrics collected over the specified interval. If the calculated traffic stress is greater than zero, an edge representing the dependency and corresponding traffic load is created. The constructed graph thus reflects both the dependency topology and associated traffic stress, enabling accurate evaluation of scheduling decisions.
Algorithm 1 Build Call-graph with dependency topology and associated traffic stress
[…]
2:Output: A directed call-graph $G$ with traffic edges
3:Initialize $G$ as an empty directed graph
[…]
7:         if $src\neq dst$ then
[…]
11:              end if
12:         end if
13:     end for
14:end for
15:return $G$
(a) Packet egress with classful disciplines.
(b) Classful qdisc for delay emulation.
(c) Classful qdisc for bandwidth shaping.
Figure 6. (a) The process of enqueueing packets and dequeuing to different destinations depending on classful disciplines. (b) Packets classified for destination-specific delay emulation. (c) Packets classified for bandwidth shaping towards different destinations.
6. Design of Networking Dynamics Manager
The Networking Dynamics Manager in the proposed framework iDynamics consists of two main components: the Emulator and Measurer. In cloud-edge computing, accurately emulating dynamic network conditions such as latency variations and bandwidth fluctuations is critical for evaluating scheduling policies against SLA targets. However, performing these evaluations directly in production environments is challenging and inefficient. iDynamics addresses this issue by providing a controllable networking dynamics emulator that accurately replicates delays and bandwidths. Moreover, precise and efficient measurement of these emulated dynamics is essential for reliable evaluation.
6.1. Linux Traffic Control
tc: tc(traffic control) is a powerful Linux utility for managing and shaping network traffic at the Linux kernel level, enabling precise control over packet transmission in complex networking scenarios. It allows fine-grained control of the flow of network packets, manages bandwidth, and prioritizes traffic on network interfaces. As part of the iproute2 suite333iproute2 suite is a modern collection of utilities designed for managing and configuring networking in Linux kernel, including routing, network interfaces, tunnels, traffic control, and network-related device drivers, tc allows users to configure queuing disciplines (qdiscs), traffic classes, and filters to emulate various networking conditions, such as bandwidth limitations, packet delays, or loss.
qdisc: qdisc(queueing discipline) defines how packets are enqueued and dequeued from a network interface. It can be used to design customized delays, manage bandwidth, prioritize traffic, and shape network traffic to meet specific rules. There are two main types of qdiscs, including classless qdiscs, which do not support classes and are simpler to configure, and classful qdiscs, which support different classes and allow for hierarchical control.
u32 filter: Filters in tc play a crucial role in directing traffic to specific classes or queues by identifying and matching packet characteristics. u32 filter is one of the most versatile and widely used filters in tc. It allows users to match arbitrary bits of a packet header, enabling highly specific rules. For example, users can match fields like source/destination IPs, port numbers, and protocol types by extracting specific bits from the header.
For the dynamic network emulations in iDynamics, classful qdiscs and u32 filter are implemented to meet the goals of changing cross-node communication delays and bandwidths in the cloud-edge continuum. Moreover, after injecting the dynamic delays or shaping available bandwidths, Cross-node Measurer is designed to measure these delays and bandwidths in the current cloud-edge continuum because the accurate and efficient measurement of these dynamics is crucial for designing and evaluating networking-related scheduling policies.
6.2. Emulator: A classful qdisc design
Emulation for Customized Delays: Emulating diverse cross‐node delays among cluster nodes is crucial for evaluating the robustness of various scheduling policies in the cloud‐edge continuum. However, implementing customized communication delays from a single source node to multiple destination nodes poses significant challenges. To the best of our knowledge, as discussed in related work, none of the existing studies has proposed an efficient method (i.e., one that is both simple and rapid) for injecting tailored communication delays in a controllable manner. In practical cloud‐edge continuum environments, these approaches generally exhibit two main limitations: (1) the use of uniform communication delays from a single source node to all destination nodes prevents differentiation between node pairs, as delays from the source to all other nodes are identical; and (2) other networking services that do not involve the correlated nodes suffer degradation because the injected delays affect all outgoing traffic.
To address these limitations, we propose a customized delay injection scheme that classifies packets using filters to distinguish egress packets and direct them based on their IP destinations (see Figure 6(b)). Additionally, we reserve an extra channel for default packet transmission without injected delays, ensuring that the performance of other services remains unaffected, for example, the packets from/to the master node and outside internet services like Google. We implemented this scheme using the tc networking tool and the Hierarchical Token Bucket (htb).
Furthermore, we employ the technique illustrated in Figure 6 to design an algorithm for emulating customized cross-node delays, as presented in Algorithm 2. To simulate cross-node delays more realistically, we incorporate several factors: the base latency bl (representing the ideal minimal latency); the maximum additional latency mal; the distance factor df (accounting for latency due to physical separation); and the congestion factor cf (which simulates network uncertainties induced by traffic congestion). Here, $\textsc{RandUni}(0,mal)$ denotes the additional delay generated according to a random uniform distribution between 00 and $mal$. Thus, the emulated communication delay from node $i$ to node $j$ in a cluster of $N$ cloud-edge nodes is given by the following equation:
(2) $\text{delay}^{j}_{i}=\left\lfloor\left(bl+\textsc{RandUni}(0,mal)\times\frac{\lvert i-j\rvert}{N}\right)\times cf\right\rfloor$
Algorithm 2 Generate Cross-Node Delay Matrix
1:Input: Number of nodes $N$, Base latency $bl$, Maximum additional latency $mal$
2:Output: Cross-node delay matrix $delay\_matrix$ of size $N\times N$
3:Initialize $delay\_matrix\leftarrow$ $[0]_{N\times N}$triangleright all set to 00
4:for each node $i$ in $[0,N-1]$ do
5:     for each node $j$ in $[0,N-1]$ do
6:         if $i==j$ then
7:              $delay\_matrix[i][j]\leftarrow 0$triangleright avoid self-delay
8:         else
9:              /*additional latency*/
10:              $al\leftarrow$ RandUniform($0,mal$)
11:              /*distance factor*/
12:              $df\leftarrow\frac{|i-j|}{N}$
13:              $emu\_latency\leftarrow bl+(al\times df)$
14:              /*congestion factor*/
15:              $cf\leftarrow$ RandUniform($0.5,1.5$)
16:              /* convert to integer */
17:              $delay\_matrix[i][j]\leftarrow\lfloor emu\_latency\times cf\rfloor$
18:         end if
19:     end for
20:end for
21:return $delay\_matrix$triangleright dynamic cloud-edge delays
Emulation for Customized Bandwidths: Shaping customized available bandwidths between different cloud-edge node pairs is also crucial for creating dynamic networking conditions, thereby enhancing the generalizability of evaluated scheduling policies. Prior research—such as the work presented in (Gouveia et al., 2020) and (Liechti et al., 2019)—has demonstrated the feasibility of emulating dynamic network conditions, including bandwidth variations, using decentralized emulation techniques. However, these approaches typically apply a global or node-level configuration that does not allow fine-grained, per-destination control.
In contrast, our proposed emulation scheme leverages the classful qdisc with unique traffic classes and utilizes the u32 filter to match different IP destinations. As illustrated in Figure 6(c), this approach enables the emulation of distinct bandwidth settings for different destination nodes, even when originating from the same source cluster node. This fine-grained control is essential for accurately emulating heterogeneous network conditions in cloud-edge environments.
6.3. Measurer: Distributed Agent Design
In practical cloud-edge computing environments, fluctuating network conditions—including varying delays and bandwidths between nodes—significantly impact microservice performance and SLA compliance. High latency between cluster nodes negatively affects microservice communication, while dynamically changing bandwidth can lead to congestion, increased packet loss, and potential SLA violations. To accurately capture these dynamic cross-node conditions, we introduce the Measurer, a distributed agent-based measurement module within the Networking Dynamics Manager component of iDynamics (Figure 4).
Figure 7. Design of Measurer for delay and bandwidth measurement through distributed agents across cloud-edge nodes
Design Overview: The Measurer uses a unified approach for capturing cross-node delays and bandwidths through distributed measuring agents. Although both delay and bandwidth measurements follow a similar structure, we illustrate the design using delay measurement for clarity. The module consists of a centralized information processing unit and distributed lightweight agents. The centralized processing unit maintains minimal connections with agents, aggregating the collected data efficiently.
The distributed agents run as lightweight containers deployed across cluster nodes, dedicated to measuring and reporting network metrics. To ensure robustness and adaptability during cluster scaling, we implement an automatic scaling mechanism for these agents. When new nodes join the cluster, corresponding agents are automatically instantiated, ensuring consistent and accurate measurements. Conversely, when nodes are removed, their associated agents are gracefully terminated.
Implementation Details: The centralized information processing unit is implemented as a plugin running on the master node, collecting measured metrics from distributed agents via standard TCP communications. The distributed agents are deployed as Kubernetes DaemonSet pods, ensuring each node automatically hosts a dedicated measurement pod. These agents continuously measure and report cross-node delays and bandwidths, dynamically adjusting their presence based on cluster scaling events, as depicted in Figure 7.
Overhead Analysis: Regarding the overhead of measurement in a real cloud-edge cluster, we optimized the measurement using parallel processing. The measurement tasks for delays and bandwidths are designed to run as concurrent tasks, and each of the measured results is aggregated into a result dictionary respectively. Additionally, the distributed delay measuring agents are lightweight (about 0.2 MiB) and stable (running over 6 months without any failure). Each node runs a delay-measuring agent developed from the curlimages/curl image, consuming around 0.2 MiB of memory per node. Similarly, each node also runs a bandwidth-measuring agent implemented from the networkstatic/iperf3 image, consuming around 0.84 MiB of memory on each node. As the measurement tasks are designed to run in parallel as well, the measuring speed can be adaptively tuned by increasing or decreasing the concurrency, depending on the current load levels in the cluster.
7. Design of Scheduling Policy Extender
The Scheduling Policy Extender provides a customizable policy interface and a utility function module designed to facilitate the rapid prototyping and evaluation of scheduling policies. By extending the provided interfaces and leveraging utility functions, users can efficiently develop tailored scheduling strategies that meet the diverse performance, scalability, and reliability requirements of microservice-based applications in cloud-edge environments.
Figure 8. Overview of the Scheduling Policy Extender, highlighting extensible classes, utility functions, and example policies.
7.1. Policy Customization Interface
The policy interface is defined through an abstract class, AbstractSchedulingPolicy, which establishes the fundamental structure required for creating scheduling strategies. This abstract class includes predefined attributes and abstract methods essential for scheduling decisions. Scheduling decisions are encapsulated into combined NodeInfo and PodInfo objects. The NodeInfo class includes node-specific attributes such as resource capacities (e.g., CPU, memory) and network characteristics (e.g., latency, bandwidth). The PodInfo class contains essential pod details, including resource requests, limits, and SLA requirements (e.g., throughput, response time), and is extensible to accommodate additional metrics.
The AbstractSchedulingPolicy also provides customizable internal methods such as single-pod scheduling (for single-service optimization), batch-pod scheduling (for multi-service optimization), and an update-metrics method triggered during re-scheduling events. By overriding these abstract methods, researchers and developers can incorporate sophisticated decision-making algorithms—ranging from heuristics to advanced machine learning techniques—while seamlessly integrating their logic with the underlying scheduling framework.
7.2. Utility Function Module
The utility function module complements the policy interface, providing ready-to-use functions that simplify and accelerate the extension of custom policies. At the node level, it includes functions for collecting real-time node resource usage and network metrics, transforming them into structured objects compatible with scheduling algorithms. At the pod level, utility functions assist in managing pod resource specifications, SLA requirements, and metric unit conversions. Furthermore, additional helper functions utilize monitoring tools such as Prometheus to gather comprehensive cluster-level metrics and network conditions. By leveraging these pre-built utility functions, users can rapidly and effectively develop sophisticated scheduling policies tailored to specific operational contexts.
7.3. Examples of Customized Policies
To illustrate the flexibility of the provided policy interface and utility functions, we present several examples of custom scheduling policies addressing distinct cloud-edge dynamics:
7.3.1. Policy 1: Call-graph–Aware
Microservice applications often exhibit dynamic call-graphs triggered by different request types, resulting in varying communication intensities between service pairs. This policy strategically co-locates microservices with high mutual communication on the same node or nodes with superior interconnectivity, thus minimizing cross-node traffic. To avoid potential node overload (e.g., excessive microservice instances in one node), stricter node capacity checks can be incorporated into the policy.
7.3.2. Policy 2: Latency-Aware
In cloud-edge deployments where inter-node latency fluctuates significantly, reducing end-to-end latency becomes essential for latency-sensitive microservices. This policy places latency-critical services or closely interacting service pairs on nodes with minimal cross-node delays, including co-locating them whenever possible to maximize performance.
7.3.3. Policy 3: Bandwidth-Aware
Certain microservices generate substantial data transfers, making them bandwidth-sensitive. In scenarios with dynamic or constrained bandwidth conditions, this policy assigns high-bandwidth-demanding microservices to nodes offering the best available bandwidth or co-locates them to reduce inter-node data movement, thus optimizing throughput and performance.
7.3.4. Policy 4: Hybrid-dynamics-Aware
Considering the combined dynamics of varying call-graphs and fluctuating cross-node delays, this policy formulates the microservice scheduling as a Service-Node Mapping Problem. It strategically maps microservices to nodes by jointly optimizing the communication cost, which depends on both inter-service traffic volume and inter-node communication delays. Specifically, microservices with high interaction volumes are placed on nodes with low mutual communication delays. This integrated approach minimizes overall communication latency and maximizes performance under realistic, dynamic cloud-edge conditions.
8. Performance Evaluation
Table 1. Injected versus measured cross-node communication delays (ms). Each cell shows the injected delay (left) and the corresponding measured actual delay (right). The results demonstrate successful injection of varied delays from source nodes to multiple destinations, along with reserved channels without injected delays (i.e., Master node and google.com).
Source Node Destinations
k8s-worker-1 k8s-worker-2 k8s-worker-3 k8s-worker-4 k8s-worker-5 k8s-worker-6 k8s-worker-7 k8s-worker-8 k8s-worker-9 Master Node google.com
k8s-worker-1 - / - 3.00 / 3.21 8.00 / 8.51 10.00 / 11.01 14.00 / 14.21 6.00 / 6.73 27.00 / 28.98 13.00 / 13.89 21.00 / 22.31 - / 0.64 - / 14.10
k8s-worker-2 8.00 / 8.22 - / - 4.00 / 4.52 13.00 / 14.02 14.00 / 15.21 18.00 / 19.02 38.00 / 38.22 31.00 / 32.13 29.00 / 1.02 - / 0.89 - / 14.20
k8s-worker-3 4.00 / 4.03 12.00 / 12.37 - / - 8.00 / 9.05 18.00 / 18.66 11.00 / 11.89 4.00 / 5.75 12.00 / 13.41 15.00 / 16.89 - / 0.43 - / 14.20
k8s-worker-4 17.00 / 17.04 15.00 / 15.11 7.00 / 7.21 - / - 5.00 / 5.87 6.00 / 6.79 22.00 / 23.15 13.00 / 13.76 25.00 / 26.82 - / 0.51 - / 14.30
k8s-worker-5 20.00 / 21.02 12.00 / 12.23 11.00 / 11.68 10.00 / 11.27 - / - 9.00 / 9.73 7.00 / 7.13 4.00 / 4.11 9.00 / 9.23 - / 0.45 - / 14.10
k8s-worker-6 17.00 / 17.12 26.00 / 26.06 18.00 / 18.87 16.00 / 17.08 6.00 / 6.11 - / - 5.00 / 5.71 10.00 / 10.19 5.00 / 5.86 - / 0.53 - / 14.20
k8s-worker-7 20.00 / 20.96 10.00 / 10.52 10.00 / 10.91 9.00 / 9.86 11.00 / 12.92 5.00 / 5.67 - / - 5.00 / 5.71 9.00 / 9.39 - / 0.45 - / 14.10
k8s-worker-8 21.00 / 21.11 25.00 / 25.72 4.00 / 4.53 10.00 / 10.71 12.00 / 12.73 15.00 / 16.02 10.00 / 10.13 - / - 6.00 / 7.08 - / 0.34 - / 14.10
k8s-worker-9 36.00 / 36.93 22.00 / 22.70 40.00 / 40.39 9.00 / 10.08 25.00 / 25.69 8.00 / 8.16 7.00 / 7.23 6.00 / 6.18 - / - - / 0.30 - / 14.30
Table 2. Saturated versus measured cross-node bandwidths (Mbits/sec). Each cell shows the saturated bandwidth first, followed by the actual measured bandwidth, which is emulated and measured by our proposed components in iDynamics.
Source Node Destinations
k8s-worker-1 k8s-worker-2 k8s-worker-3 k8s-worker-4 k8s-worker-5 k8s-worker-6 k8s-worker-7 k8s-worker-8 k8s-worker-9
k8s-worker-1 - / - 594 / 608 659 / 665 437 / 452 345 / 359 550 / 560 277 / 300 755 / 754 659 / 667
k8s-worker-2 251 / 267 - / - 512 / 517 270 / 294 432 / 446 404 / 419 274 / 274 625 / 620 386 / 406
k8s-worker-3 721 / 685 512 / 507 - / - 485 / 485 300 / 305 439 / 439 340 / 342 234 / 245 751 / 711
k8s-worker-4 427 / 434 772 / 769 201 / 206 - / - 692 / 682 238 / 242 427 / 449 594 / 598 252 / 253
k8s-worker-5 385 / 376 398 / 397 467 / 467 622 / 606 - / - 288 / 302 501 / 492 502 / 484 683 / 582
k8s-worker-6 675 / 679 779 / 565 247 / 269 229 / 232 484 / 500 - / - 429 / 422 230 / 243 698 / 553
k8s-worker-7 555 / 569 467 / 477 240 / 259 534 / 545 580 / 589 202 / 221 - / - 449 / 466 737 / 722
k8s-worker-8 313 / 330 361 / 336 427 / 442 628 / 548 419 / 434 707 / 675 564 / 528 - 393 / 411
k8s-worker-9 605 / 619 400 / 389 346 / 362 693 / 562 372 / 389 748 / 710 783 / 736 566 / 576 - / -
8.1. Cloud-edge Testbed Setup
Setup: We implemented and validated the proposed iDynamics framework using Kubernetes (open-source container orchestration system, 2024), the de facto container orchestration platform. The experimental cloud-edge cluster comprised one master node and nine worker nodes without preset anti-colocation rules such as node taints or pod affinity. The master node features 32 CPU cores (AMD EPYC 7763, x86_64 architecture), 32 GiB RAM, and 16 Gbps network bandwidth. Each worker node contains 4 CPU cores from the same AMD series, 32 GiB RAM, and 16 Gbps network bandwidth. The software stack includes Kubernetes v1.27.4, Calico v3.26.1 (CNI plugin), Istio v1.20.3 (service mesh), and CRI-O v1.27.1 as the container runtime. All nodes operate on Ubuntu 22.04.2 LTS with Linux kernel 5.15.0.
The cluster nodes are virtual machines hosted on the university’s dedicated research cloud, providing ultra-low communication delays between nodes (typically between 0.2 and 1 ms, verified by ICMP tests). To emulate realistic and dynamic cloud-edge networking conditions, we inject customized cross-node delays and bandwidth constraints into the cluster nodes. These conditions are periodically updated to evaluate the adaptability of iDynamics under varying network dynamics.
Workload Generation: We adopted Social Network from DeathStarBench(Gan et al., 2019) as the microservice application workload and wrk2 (wrk2: An HTTP benchmarking tool based on wrk, 2024) as the requests generator to validate iDynamics and the proposed sample scheduling policies. Social Network benchmark emulates a simplified social media platform similar to popular social networking services. It is structured to replicate the intricate interactions and communication patterns in such types of applications. The benchmark comprises 27 different microservices that collectively offer request types such as composing a post, writing a user timeline, and writing a home timeline, which would trigger different call-graph topologies. wrk2 is capable of generating different types and proportions of workload requests for performance testing and measuring how well the cloud applications can respond to varying traffic. By combining the advantages of Social Network application and wrk2 benchmarking tool, various scenarios of workload generation can be achieved, thus facilitating the generation of call-graph dynamics and evaluation of different scheduling policies.
8.2. Effectiveness and Validation of Networking Dynamics Manager
Effectiveness: Precisely injecting and measuring diverse communication delays between nodes is challenging. We validated our delay injection method (depicted in Figure 6(b)) by injecting varied delays from a source node to multiple destinations and measuring the actual communication delays. Results in Table 1 demonstrate high accuracy, typically within 1 ms of the target delay, minor variations attributed to inherent randomness in cloud networking stacks. Additionally, to demonstrate that other traffic is not influenced, we tested the communication delays from all worker nodes to other destinations. One destination is the master node in the same cluster, and the other is the Google host (google.com), showing average delays of 0.50 $ms$ and 14.20 $ms$, respectively.
Similarly, bandwidth constraints were effectively enforced using our bandwidth shaping approach (illustrated in Figure 6(c)), as shown in Table 2. The measured bandwidth closely aligns with target values, confirming the efficacy of the method.
[…]
8.3. Case Studies for Policy Evaluations
Table 3. Performance comparison between Kubernetes default scheduling and Policy 1 (Call-graph–Aware) under varying workloads and call-graph dynamics. Metrics include average (Avg), 99th percentile (p99), and standard deviation (Stdev) of response time (ms). SLA violations (¿150 ms) are highlighted in red, and SLA-compliant improvements are highlighted in green.
Policy Running time 0–8 min (Call-graph 1) 8–16 min (Call-graph 2) 16–24 min (Call-graph 3) 24–32 min (Call-graph 4, mixed)
QPS (req/sec) 30 10 50 70 30 10 50 70 30 10 50 70 30 10 50 70
K8s Policy Avg 13.27 14.36 13.22 13.22 168.6 171.8 171.2 174.2 57.33 56.06 49.92 46.11 40.67 44.15 39.31 40.15
p99 20.24 38.91 18.08 18.64 299.1 243.8 315.6 326.9 197.8 203.1 197.3 187.5 207.7 259.7 203.5 203.52
Stdev 2.46 4.61 2.51 1.88 24.87 20.17 33.17 39.46 48.3 45.73 38.64 31.42 51.43 60.34 49.63 50.49
Policy 1 Avg 13.89 14.84 13.51 13.33 168.88 127.6 117.26 118.4 71.32 70.87 68.4 61.83 34.92 39.37 35.2 35.69
p99 23.85 42.72 17.38 20.00 300.8 594.4 210.7 218.75 309.76 314.4 310.0 305.4 202.2 251.3 183.4 194.3
Stdev 3.26 5.26 2.10 2.13 25.27 92.80 19.12 22.33 68.42 68.38 63.62 49.84 43.76 51.00 43.75 43.88
8.3.1. Policy 1: Call-graph Scheduling Policy Evaluation
Figure 9. Average response time comparision between K8s-scheduling and Policy1 (Call-graph-Aware scheduling) under sustained and varying workloads.
In this case study, we evaluate the effectiveness of the Call-graph–Aware scheduling policy, which strategically places microservices with intensive inter-service communication onto nodes with high interconnectivity, reducing cross-node traffic.
Dynamic Workloads (QPS and Call-graphs): Experiments were conducted under dynamic workloads generated by the wrk2 tool, reflecting realistic call-graph dynamics. We utilized four distinct call-graph topologies, each tested with varying queries per second (QPS), thus creating scenarios with changing call-graph topologies and traffic intensities.
Discussion: Figure 9 and Table 3 clearly demonstrate Policy 1’s capability to rapidly mitigate SLA violations compared to the default Kubernetes scheduler. Upon SLA violation at the 10-minute point, Policy 1 effectively reduced the average response time from approximately 170 ms to about 120 ms, maintaining SLA compliance during heavy workloads. In contrast, Kubernetes default scheduling continued violating the SLA until the workload intensity decreased.
8.3.2. Policy 4: Hybrid-dynamics–Aware Policy Evaluation
Figure 10. Average response time comparision between Policy1 (Call-graph-Aware scheduling and Policy 4 (Hybrid-dynamics-Aware) under sustained workloads and changing networking conditions.
The Hybrid-dynamics–Aware policy integrates considerations of both call-graph dynamics and fluctuating network conditions (cross-node latency). It formulates microservice scheduling as a Service-Node Mapping Problem, aiming to jointly minimize communication costs arising from inter-service traffic volumes and network latencies.
Dynamic Workloads (QPS and Networking Conditions): The performance of Policy 4 was evaluated under sustained and variable QPS and randomized cross-node latency scenarios, emphasizing the policy’s adaptability to multiple dynamics simultaneously.
Discussion: As depicted in Figure 10, both policies maintained response times below the SLA target (¡ 300 ms). However, Policy 4 (Hybrid-dynamics–Aware) generally outperformed Policy 1 throughout the 40-minute evaluation. This confirms the Hybrid-dynamics–Aware policy’s enhanced ability to adapt dynamically and efficiently to combined network and QPS fluctuations, validating its suitability for complex and dynamic cloud-edge environments.
9. Conclusions and Future Work
In this paper, we presented iDynamics, a novel evaluation framework designed to systematically assess microservice scheduling policies under various dynamics in cloud-edge computing environments, particularly addressing call-graph dynamics and cross-node networking variability. The modular design of iDynamics enables realistic modelling and emulation of these dynamic conditions, along with providing extensible interfaces for users to develop and evaluate customized scheduling policies effectively.
For future work, we plan to extend the current framework by incorporating larger-scale clusters and heterogeneous resource provisioning capabilities. Additionally, we aim to explore and evaluate the performance of edge-intelligence workloads within a dynamically evolving cloud-edge continuum.